Solve: A certain light bulb containing argon has a pressure of 1.20 atm at 18°C. If it

will be heated to 85°C at constant volume, what will be the resulting pressure? 


1
Expert's answer
2022-05-02T14:54:04-0400

P1 = 1.20 atm

T1 = 18°C + 273.15 = 291.15 K

V1 = V2 = constant

P2 = ???

T2 = 85°C + 273.15 = 358.15 K


Solution:

Since the volume and amount of gas remain unchanged, Gay-Lussac's law can be used.

Gay-Lussac's law can be expressed as: P1/T1 = P2/T2

P1T2 = P2T1

To find the resulting pressure, solve the equation for P2:

P2 = P1T2 / T1

P2 = (1.20 atm × 358.15 K) / (291.15 K) = 1.476 atm = 1.48 atm

P2 = 1.48 atm


Answer: The resulting pressure will be 1.48 atm